# ReflectDiffu: Reflect between Emotion-intent Contagion and Mimicry for Empathetic Response Generation via a RL-Diffusion Framework

### v2

@@ Abstract
Empathetic response generation necessitates the integration of emotional and intentional dynamics to foster meaningful interactions. Existing research either neglects the intricate interplay between emotion and intent, leading to suboptimal controllability of empathy, or resorts to large language models (LLMs), which incur significant computational overhead. In this paper, we introduce ReflectDiffu, a lightweight and comprehensive framework for empathetic response generation. This framework incorporates emotion contagion to augment emotional expressiveness and employs an emotion-reasoning mask to pinpoint critical emotional elements. Additionally, it integrates intent mimicry within reinforcement learning for refinement during diffusion. By harnessing an intent twice reflect the mechanism of Exploring-Sampling-Correcting, ReflectDiffu adeptly translates emotional decision-making into precise intent actions, thereby addressing empathetic response misalignments stemming from emotional misrecognition. Through reflection, the framework maps emotional states to intents, markedly enhancing both response empathy and flexibility. Comprehensive experiments reveal that ReflectDiffu outperforms existing models regarding relevance, controllability, and informativeness, achieving state-of-the-art results in both automatic and human evaluations.
@@
Jiahao Yuan1, Zixiang Di1, Zhiqing Cui2, Guisong Yang3*, Usman Naseem4* 1East China Normal University 2Nanjing University of Information Science and Technology 3University of Shanghai for Science and Technology 4Macquarie University {51275900024,51265901113}@stu.ecnu.edu.cn, 202383090003@nuist.edu.cn, gsyang@usst.edu.cn, usman.naseem@mq.edu.au
@@ 1 Introduction
Recent advances in large language models (LLMs) Dubey et al. (2024); Yang et al. (2024) have promoted several empathetic dialogue models utilizing multiple-stage Chain-of-Thought (COT) Chen et al. (2024); Hu et al. (2024) with fine-tuning Zhang et al. (2020); Cai et al. (2024). However, their unstable performance Lu et al. (2022); Xie et al. (2023) and reliance on external knowledge and high training costs Kaplan et al. (2020) complicate practical implementation. Consequently, current research focuses on enhancing small-scale empathetic models through empathy mechanisms Majumder et al. (2020); Gao et al. (2023); Zhou et al. (2023); Wang et al. (2024) as a more lightweight, practical alternative to LLMs. In summary, lightweight empathetic models encounter three major limitations: (1) They primarily rely on supplementary knowledge signals Gao et al. (2023) rather than underlying psychological mechanisms, which impedes controllability and empathetic capability. (2) They often overlook the internal mechanisms behind emotional causes, emotions, and intents, which rely heavily on external knowledge or pre-trained annotators Bi et al. (2023); Chen et al. (2024), resulting in hard-coded enhancements rather than genuine understanding and iterative correction, thus impacting empathy, diversity and flexibility. (3) There is a shortage of multi-task datasets for emotion reason masking, intent prediction, and empathetic dialogue. Most models rely on supplementary datasets for auxiliary tasks Li et al. (2024); Bi et al. (2023), which does not guarantee that the advantages of multi-task training are fully realized or effectively aligned.
@@
  • We utilize LLMs to expand annotation on EmpatheticDialogue Rashkin et al. (2018) and create a multi-task dataset for emotional reason masking, emotion prediction, intent prediction, and empathetic dialogue generation. Our dataset is available [^1] https://anonymous.4open.science/r/ReflectDiffu-4C12.
@@ 2.2 Generative Model for Dialogue Generation
Subsequently, Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. (2020) stands out for generating high-quality samples via iterative denoising. Hoogeboom et al. (2021) and Austin et al. (2021) paved the way for character-level text generation with diffusion. Li et al. (2022b) uses an embedding and rounding strategy and additional classifiers for controllable text generation. Gong et al. (2022) introduces a classifier-free diffusion model for dialogue generation. Bi et al. (2023) incorporated multi-grained control signals but their multi-stage pre-training approaches increase computational costs and the difficulty of practical implementation.
@@ 3.2 Multi-task Emotion-Contagion Encoder
Inspired by Lin et al. (2019); Majumder et al. (2020); Chen et al. (2022b), we put forward two-expert models C-Experts: $\mathbbm{M}{pos}$ for positive emotions and $\mathbbm{M}{neg}$ for negative emotions to enhance emotion recognization by exploiting each model’s proficiency. Neutral emotions are addressed via a voting mechanism between experts, yielding the candidate emotion probability distribution $p$ as follows::
$p=\begin{cases}\texttt{softmax}(W_{\textit{neg}}\mathbf{E_{\textit{emo}}}Q)&\textit{if }v=\textit{n}_{\textit{neg}}\\ \text{softmax}(W_{\textit{pos}}\mathbf{E_{\textit{emo}}}Q)&\textit{if }v=\textit{n}_{\textit{pos}}\\ Voting(W_{\text{neg}}\mathbf{E_{\texttt{emo}}}Q),W_{\text{pos}}\mathbf{E_{\texttt{emo}}}Q)))&\text{if }v=\text{n}_{\text{neu}}\\ \end{cases}$ (4)
@@ 3.3.2 Response Decoding:
$=P_{\text{Gen}}(\texttt{TRS}_{\texttt{dec}}(E^{C}(T_{R<t}),\textit{Emo}_{\textit{fused}}),$ (11)
@@ 4 Experiments Settings
BLEU-1 $\uparrow$ BLEU-2 $\uparrow$ BLEU-3 $\uparrow$ BLEU-4 $\uparrow$ BARTScore $\uparrow$ Accemo $\uparrow$ Accintent $\uparrow$ PPL $\downarrow$ Dist-1 $\uparrow$ Dist-2 $\uparrow$
MTRS Rashkin et al. (2018) 17.87 8.51 4.36 2.61 0.5173 32.96 - 37.98 0.40 1.57
MOEL Lin et al. (2019) 18.02 8.67 4.35 2.73 0.5166 31.02 - 36.81 0.43 1.76
MIME Majumder et al. (2020) 19.82 8.86 4.43 2.77 0.5182 30.26 - 36.93 0.51 1.92
EmpDG Li et al. (2020) 19.12 8.91 4.89 2.85 0.5171 32.90 - 37.55 0.49 1.65
KEMP Li et al. (2022a) 17.92 8.54 4.38 2.71 0.5232 36.40 - 36.59 0.66 2.43
CASE Zhou et al. (2023) 19.66 8.95 4.92 2.90 0.5336 38.96 - 35.97 0.70 2.66
CAB Gao et al. (2023) 20.23 9.39 4.96 3.01 0.5392 40.52 - 35.06 0.89 2.95
Qwen2-7B + COTYang et al. (2024) 23.31 11.21 5.20 3.45 0.5447 23.10 41.61 25.45 0.87 3.87
llama-3.1-8B + COTDubey et al. (2024) 23.38 11.29 5.25 3.47 0.5480 21.15 32.02 24.92 0.92 4.13
@@
ReflectDiffu w/o ERA 22.59 10.66 5.02 3.28 0.5520 42.37 78.68 24.78 0.95 4.27
ReflectDiffu w/o C-Experts 23.13 11.05 5.06 3.31 0.5619 39.44 77.44 24.82 0.91 4.03
ReflectDiffu w/o Intent twice 20.91 9.86 4.87 3.16 0.5436 44.56 66.44 29.25 0.85 3.97
ReflectDiffu w/o EMU 21.95 10.05 4.96 3.22 0.5490 48.35 79.24 27.45 0.69 2.96
Table 2: Results of automatic evaluations and ablation study.
@@ 4.2 Baselines
In our experiments, we compare ReflectDiffu with both classic and recent state-of-the-art (SOTA) benchmarks including MTRS Rashkin et al. (2018) , MOEL Lin et al. (2019) , MIME Majumder et al. (2020) , EmpDG Li et al. (2020) , KEMP Li et al. (2022a) , CASE Zhou et al. (2023) , and CAB Gao et al. (2023). Additionally, we incorporate a comparative analysis with QWen2-7B Yang et al. (2024) and Llama-3.1-8BDubey et al. (2024), two prominent generative language modelsLaskar et al. (2024). More details about baselines are shown in Appendix A.1.
@@ 4.4 Evaluation Metrics
For human evaluation, we conduct A/B testing on empathy, relevance and fluency with three recruited annotators and a supervisory LLM to resolve disagreements (see Appendix A.2 for details).
@@ 5 Results and Discussion
As shown in table 2, our model, ReflectDiffu, outperforms all baseline models and significantly enhances all metrics. Compared with empathy-specific modelsRashkin et al. (2018); Lin et al. (2019); Majumder et al. (2020); Li et al. (2020, 2022a); Zhou et al. (2023); Gao et al. (2023), which mainly explore the connection between emotion states and empathetic contexts but ignore the internal mechanisms of emotional causes, emotions, and intents and only rely on inferred external knowledge, resulting in suboptimal empathetic controllability (low emotion accuracy $Acc_{\textit{emo}}$), weak similarity and coherence with the empathetic ground truth (indicated by low BLEU-n and $\textit{BART}_{\textit{Score}}$), and a lack of diversity (implied by low Distinct-1 and Distinct-2). In contrast, ReflectDiffu exhibits remarkable superiority, exceeding the best baseline, CAB, approximately in BLEU-1, BLEU-2, BLEU-3, BLEU-4, $\textit{BART}_{\textit{Score}}$, $\textit{Acc}_{\textit{emo}}$ by 16.6% , 20%, 8.1%, 20.3%, 4.6% and 20.3% respectively for Emotion-Contagion Encoder to enhance semantic understanding and achieve 80.32% intent accuracy for its intent twice mechanism. Moreover, ReflectDiffu shows improvements of approximately 30.1% in $PPL$, 10.1% in Distinct-2, and 47.4% in Distinct-2 compared with CAB for Diffusion within intent guidance.
@@ 6 Conclusion
We discuss the primary limitation of this work as follows: The integration of Denoising Diffusion Probabilistic Models (DDPMs) and reinforcement learning mechanisms has augmented the computational requirements for training, presenting challenges for deployment in resource-constrained settings or on devices with limited capabilities. To alleviate this limitation, we have adopted reparameterization and multi-task techniques for optimization. As a result, the overall training time is notably shorter than that of multi-stage LLM Chen et al. (2024); Yang et al. (2024); Dubey et al. (2024) while achieving state-of-the-art outcomes.
@@ A.1 Baselines
  • QWen2-7B + COT: We fine-tune QWen2-7BYang et al. (2024), and then employ Chain-of-Thought (COT) to infer emotion, intent, and generate empathetic responses for improved empathy.
@@ B.2 Case Study in Misclassification
We deliberately selected a ReflectDiffu emotion recognition error case to validate the effectiveness of our reflection mechanism. Table 5 compares responses from various models, including MOEL, MIME, EmpDG, KEMP, CASE, CAB, Qwen2-7B+COT,llama3.1-8B+COT and ReflectDiffu, to a user’s context of feeling hopeful after applying for graduate school. Initially, ReflectDiffu misidentified the emotion as "joyful" and the intent as "acknowledging." However, after employing the reflection mechanism, it correctly identified the intent as "encouraging." This demonstrates the model’s capability to correct errors and generate more empathetic responses through its reflection mechanism.
@@ C.1 Emotion Reason Annotator
$\alpha_{ik}^{jm}=\frac{\exp\left(\text{Attention}(h^{i}_{j},h^{k}_{m})\right)}{\sum_{k,m}\exp\left(\text{Attention}(h^{i}_{j},h^{k}_{m})\right)},$ (14)
@@ C.2 Definition of $L_{\text{em}}$
Appendix D Annotators Instructions for Human Evaluation.
@@
  • Evaluation Criteria: Annotators assessed responses based on three key criteria::

### v4

@@ Abstract
Empathetic response generation necessitates the integration of emotional and intentional dynamics to foster meaningful interactions. Existing research either neglects the intricate interplay between emotion and intent, leading to suboptimal controllability of empathy, or resorts to large language models (LLMs), which incur significant computational overhead. In this paper, we introduce ReflectDiffu, a lightweight and comprehensive framework for empathetic response generation. This framework incorporates emotion contagion to augment emotional expressiveness and employs an emotion-reasoning mask to pinpoint critical emotional elements. Additionally, it integrates intent mimicry within reinforcement learning for refinement during diffusion. By harnessing an intent twice reflect mechanism of Exploring-Sampling-Correcting, ReflectDiffu adeptly translates emotional decision-making into precise intent actions, thereby addressing empathetic response misalignments stemming from emotional misrecognition. Through reflection, the framework maps emotional states to intents, markedly enhancing both response empathy and flexibility. Comprehensive experiments reveal that ReflectDiffu outperforms existing models regarding relevance, controllability, and informativeness, achieving state-of-the-art results in both automatic and human evaluations.
@@
Jiahao Yuan1,2thanks: jhyuan.cs@gmail.com, Zixiang Di2, Zhiqing Cui3, Guisong Yang1thanks: Corresponding author: gsyang@usst.edu.cn, Usman Naseem4 1University of Shanghai for Science and Technology 2East China Normal University 3Nanjing University of Information Science and Technology 4Macquarie University
@@ 1 Introduction
Recent advances in large language models (LLMs) Dubey et al. (2024); Yang et al. (2024a) have promoted several empathetic dialogue models utilizing multiple-stage Chain-of-Thought (CoT) Chen et al. (2024); Hu et al. (2024) with fine-tuning Zhang et al. (2020); Cai et al. (2024). However, their unstable performance Lu et al. (2022); Xie et al. (2023); Yuan et al. (2024) and reliance on external knowledge and high training costs Kaplan et al. (2020); Yuan et al. (2024) complicate practical implementation. Consequently, current research focuses on enhancing small-scale empathetic models through empathy mechanisms Majumder et al. (2020); Gao et al. (2023); Zhou et al. (2023); Wang et al. (2024) as a more lightweight, practical alternative to LLMs. In summary, lightweight empathetic models encounter three major limitations: (1) They primarily rely on supplementary knowledge signals Gao et al. (2023) rather than underlying psychological mechanisms, which impedes controllability and empathetic capability. (2) They often overlook the internal mechanisms behind emotional causes, emotions, and intents, which rely heavily on external knowledge or pre-trained annotators Bi et al. (2023); Chen et al. (2024), resulting in hard-coded enhancements rather than genuine understanding and iterative correction, thus impacting empathy, diversity and flexibility. (3) There is a shortage of multi-task datasets for emotion reason masking, intent prediction, and empathetic dialogue. Most models rely on supplementary datasets for auxiliary tasks Li et al. (2024); Bi et al. (2023), which does not guarantee that the advantages of multi-task training are fully realized or effectively aligned.
@@ 2.2 Generative Model for Dialogue Generation
Subsequently, Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. (2020) stands out for generating high-quality samples via iterative denoising Li et al. (2025). Hoogeboom et al. (2021) and Austin et al. (2021) paved the way for character-level text generation with diffusion. Li et al. (2022b) uses an embedding and rounding strategy and additional classifiers for controllable text generation. Gong et al. (2022) introduces a classifier-free diffusion model for dialogue generation. Bi et al. (2023) incorporated multi-grained control signals but their multi-stage pre-training approaches increase computational costs and the difficulty of practical implementation.
@@ 3.2 Multi-task Emotion-Contagion Encoder
Inspired by Lin et al. (2019); Majumder et al. (2020); Chen et al. (2022b), we put forward two-expert models C-Experts: $\mathbbm{M}{pos}$ for positive emotions and $\mathbbm{M}{neg}$ for negative emotions to enhance emotion recognization by exploiting each model’s proficiency. Neutral emotions are addressed via a voting mechanism between experts, yielding the candidate emotion probability distribution $p$ as follows:
$p=\begin{cases}\texttt{softmax}(W_{\textit{neg}}\mathbf{E_{\textit{emo}}}Q)&\textit{if }v=\textit{n}_{\textit{neg}}\\ \texttt{softmax}(W_{\textit{pos}}\mathbf{E_{\textit{emo}}}Q)&\textit{if }v=\textit{n}_{\textit{pos}}\\ \texttt{Voting}(W_{\textit{neg}}\mathbf{E_{\textit{emo}}}Q),W_{\textit{pos}}\mathbf{E_{\textit{emo}}}Q)))&\text{if }v=\textit{n}_{\textit{neu}}\\ \end{cases}$ (4)
@@ 3.3.2 Response Decoding
$=P_{\text{Gen}}(\texttt{TRS}_{\texttt{dec}}(E^{C}(T_{R<t}),\textit{Emo}_{\textit{fused}}).$ (11)
@@ 4 Experiments Settings
B-1 $\uparrow$ B-2 $\uparrow$ B-3 $\uparrow$ B-4 $\uparrow$ BARTScore $\uparrow$ $\textit{Acc}_{\textit{emo}}\uparrow$ $\textit{Acc}_{\textit{Intent}}\uparrow$ PPL $\downarrow$ D-1 $\uparrow$ D-2 $\uparrow$
MTRS 17.87 8.51 4.36 2.61 0.5173 32.96 - 37.98 0.40 1.57
MOEL 18.02 8.67 4.35 2.73 0.5166 31.02 - 36.81 0.43 1.76
MIME 19.82 8.86 4.43 2.77 0.5182 30.26 - 36.93 0.51 1.92
EmpDG 19.12 8.91 4.89 2.85 0.5171 32.90 - 37.55 0.49 1.65
KEMP 17.92 8.54 4.38 2.71 0.5232 36.40 - 36.59 0.66 2.43
CASE 19.66 8.95 4.92 2.90 0.5336 38.96 - 35.97 0.70 2.66
CAB 20.23 9.39 4.96 3.01 0.5392 40.52 - 35.06 0.89 2.95
IAMM 19.51 8.74 4.86 3.32 0.5456 43.72 - 25.94 0.88 3.05
Harnessing (0-shot) 6.57 2.68 1.68 1.07 0.3881 24.40 - 230.99 1.79 14.72
Qwen2-7B+CoT 23.31 11.21 5.20 3.45 0.5447 23.10 41.61 25.45 0.87 3.87
llama-3.1-8B+CoT 23.38 11.29 5.25 3.47 0.5480 21.15 32.02 24.92 0.92 4.13
@@
w/o ERA 22.59 10.66 5.02 3.28 0.5520 42.37 78.68 24.78 0.95 4.27
w/o C-Experts 23.13 11.05 5.06 3.31 0.5619 39.44 77.44 24.82 0.91 4.03
w/o Intent twice 20.91 9.86 4.87 3.16 0.5436 44.56 66.44 29.25 0.85 3.97
w/o EMU 21.95 10.05 4.96 3.22 0.5490 48.35 79.24 27.45 0.69 2.96
Table 2: Results of automatic evaluations and ablation study. Metrics include BLEU-1 to BLEU-4 (B-1 B-4), BARTScore for Relevance; emotion and intent accuracy ($\textit{Acc}_{\textit{emo}}$, $\textit{Acc}_{\textit{Intent}}$) for Controllability; perplexity (PPL) and distinct-n (D-1, D-2) for Informativeness.
@@ 4.2 Baselines
In our experiments, we compare ReflectDiffu with both classic and recent state-of-the-art (SOTA) benchmarks including MTRS Rashkin et al. (2018) , MOEL Lin et al. (2019) , MIME Majumder et al. (2020) , EmpDG Li et al. (2020) , KEMP Li et al. (2022a) , CASE Zhou et al. (2023) , CAB Gao et al. (2023) and IAMM Yang et al. (2024b). Additionally, we incorporate a comparative analysis with Harnessing (0-shot prompting) Qian et al. (2023), as well as QWen2-7B Yang et al. (2024a) and Llama-3.1-8BDubey et al. (2024), two prominent generative language modelsLaskar et al. (2024). More details about baselines are shown in Appendix A.1.
@@ 4.4 Evaluation Metrics
For human evaluation, we conduct A/B testing on empathy, relevance, and fluency with three recruited annotators and a supervisory LLM to resolve disagreements (see Appendix A.2 for details). We compare ReflectDiffu against a selection of widely adopted baselines and recent generation models to ensure consistency and interpretability in human judgments. For clarity and consistency, models relying heavily on memory mechanisms (IAMM) or general-purpose prompting strategies (Harnessing) are excluded from this evaluation.
@@ 5 Results and Discussion
As shown in table 2, our model, ReflectDiffu, outperforms all baseline models and significantly enhances all metrics. Compared with empathy-specific modelsRashkin et al. (2018); Lin et al. (2019); Majumder et al. (2020); Li et al. (2020, 2022a); Zhou et al. (2023); Gao et al. (2023), which mainly explore the connection between emotion states and empathetic contexts but ignore the internal mechanisms of emotional causes, emotions, and intents and only rely on inferred external knowledge, resulting in suboptimal empathetic controllability (low emotion accuracy $Acc_{\textit{emo}}$), weak similarity and coherence with the empathetic ground truth (indicated by low BLEU-n and $\textit{BART}_{\textit{Score}}$), and a lack of diversity (implied by low Distinct-1 and Distinct-2). In contrast, ReflectDiffu exhibits remarkable superiority, exceeding the best baseline, CAB, approximately in BLEU-1, BLEU-2, BLEU-3, BLEU-4, $\textit{BART}_{\textit{Score}}$, $\textit{Acc}_{\textit{emo}}$ by 16.6% , 20%, 8.1%, 20.3%, 4.6% and 20.3% respectively for Emotion-Contagion Encoder to enhance semantic understanding and achieve 80.32% intent accuracy for its intent twice mechanism. Moreover, ReflectDiffu shows improvements of approximately 30.1% in $PPL$, 10.1% in Distinct-2, and 47.4% in Distinct-2 compared with CAB for Diffusion within intent guidance. Compared with Harnessing (0-shot), which performs poorly across relevance, controllability, and fluency (e.g., BLEU-1: 6.57, Accemo: 24.40, PPL: 230.99), ReflectDiffu achieves substantially higher scores while maintaining coherence and diversity, demonstrating its robustness in zero-shot empathetic dialogue. Compared to IAMM, which excels in emotion accuracy (43.72) and diversity (Distinct-2: 3.05) but lacks intent controllability and suffers from higher perplexity, ReflectDiffu achieves superior balance across all dimensions, with higher BLEU scores, better controllability (Accintent: 80.32), and lower PPL. These results highlight ReflectDiffu’s effectiveness in generating empathetic, coherent, and diverse responses.
@@
IAMM That sounds really scary and I’m glad you’re okay.
Harnessing So scary! Glad you’re okay—stay safe out there.
@@ 6 Conclusion
We discuss the primary limitation of this work as follows: The integration of Denoising Diffusion Probabilistic Models (DDPMs) and reinforcement learning mechanisms has augmented the computational requirements for training, presenting challenges for deployment in resource-constrained settings or on devices with limited capabilities. To alleviate this limitation, we have adopted reparameterization and multi-task techniques for optimization. As a result, the overall training time is notably shorter than that of multi-stage LLM Chen et al. (2024); Yang et al. (2024a); Dubey et al. (2024) while achieving state-of-the-art outcomes.
@@ A.1 Baselines
  • IAMM: Yang et al. (2024b) improves empathetic response quality by modeling internal affect memory and multi-level affective matching. It is designed to enhance emotional alignment and content diversity.
  • Harnessing (0-shot): Qian et al. (2023) leverages GPT-4o to generate empathetic responses via zero-shot setting under 30 budget tokens.
  • QWen2-7B + CoT: We fine-tune QWen2-7BYang et al. (2024a), and then employ Chain-of-Thought (CoT) to infer emotion, intent, and generate empathetic responses for improved empathy.
@@ B.1 Explanation of the hyperparameter $n$ of $\text{intent}_{\text{infer}}$
IAMM Perfect! Graduate school applications are exciting and I hope it works out for you.
Harnessing That’s great! Hope it goes well for you.
@@ B.2 Case Study in Misclassification
We deliberately selected a ReflectDiffu emotion recognition error case to validate the effectiveness of our reflection mechanism. Table 5 compares responses from various models, including MOEL, MIME, EmpDG, KEMP, CASE, CAB, IAMM, Harnessing, Qwen2-7B+CoT, llama3.1-8B+CoT and ReflectDiffu, to a user’s context of feeling hopeful after applying for graduate school. Initially, ReflectDiffu misidentified the emotion as "joyful" and the intent as "acknowledging." However, after employing the reflection mechanism, it correctly identified the intent as "encouraging." This demonstrates the model’s capability to correct errors and generate more empathetic responses through its reflection mechanism.
@@ C.1 Emotion Reason Annotator
$\alpha_{ik}^{jm}=\frac{\exp\left(\texttt{Attention}(h^{i}_{j},h^{k}_{m})\right)}{\sum_{k,m}\exp\left(\texttt{Attention}(h^{i}_{j},h^{k}_{m})\right)},$ (14)
@@ C.2 Definition of $L_{\text{em}}$
Appendix D Annotators Instructions for Human Evaluation
@@
  • Evaluation Criteria: Annotators assessed responses based on three key criteria: